# FMIP: Joint Continuous-Integer Flow for Mixed-Integer Linear Programming

### v2

@@ 1 Introduction
Despite its versatility, solving a MILP instance remains a fundamental challenge due to its NP-hard property (Bertsimas and Tsitsiklis, 1997). Consequently, exact solvers often rely on heuristics to find high-quality solutions in a practical amount of time. This has motivated a surge of interest in using deep learning to predict powerful heuristics that can warm-start and guide the downstream solvers (Nair et al., 2020; Han et al., 2023; Liu et al., 2025; Chen et al., 2025). Among them, generative modeling has emerged as a particularly promising approach. Modern techniques like diffusion models and flow matching reframe the difficult one-step task of predicting a complete solution into an iterative refinement process (Sun and Yang, 2023; Feng et al., 2024; Zeng et al., 2024). By learning to transform a simple noise distribution into a distribution of high-quality solutions, these models can effectively navigate the complex combinatorial search space, breaking down the challenge of satisfying intricate constraints into a series of more manageable steps.
@@
To this end, we propose Joint Continuous-Integer Flow for Mixed-Integer Linear Programming (FMIP), which is the first generative framework that models the joint distribution of both integer and continuous variables for MILP solutions. FMIP leverages a conditional flow matching process to progressively generate a complete solution, fully capturing the interdependence between all decision variables. This joint modeling approach unlocks a holistic guidance mechanism that can steer the generation process using complete, instance-wise feedback from both the objective function and constraint violations. FMIP is fully compatible with arbitrary backbone networks and downstream solvers, showing 41.34% relative improvement on average across eight standard MILP benchmarks over SOTA baselines. Our contributions are summarized as follows:
@@ 2 Related Works
Machine learning is widely used to accelerate Mixed-Integer Linear Programming (MILP) solvers. One major line of research focuses on enhancing the internal components of the Branch-and-Bound algorithm, such as learning policies for variable branching (Gasse et al., 2019a; Khalil et al., 2016; Gasse et al., 2019b; Gupta et al., 2020; Zarpellon et al., 2021; Gupta et al., 2022; Scavuzzo et al., 2022; Lin et al., 2024; Zhang et al., 2024) or cutting plane selection (Tang et al., 2020; Huang et al., 2022). Another closely related direction, which our work belongs to, aims to predict high-quality heuristic solutions to warm-start the optimization process, thereby reducing the search space and speeding up convergence (Nair et al., 2020; Han et al., 2023; Huang et al., 2024b; Liu et al., 2025; Chen et al., 2025; Zeng et al., 2024).
Most existing solution-prediction methods are discriminative in nature, employing models like Graph Neural Networks (GNNs) to predict the variable assignments in a single forward pass (Huang et al., 2024b; Liu et al., 2025; Hu et al., 2024; Huang et al., 2024a; Ye et al., 2024). However, predicting a complete, feasible, and high-quality solution in one step is exceptionally challenging due to the complex combinatorial structure of MILPs. This difficulty has motivated the use of generative models, such as diffusion models and flow matching, for MILP heuristics (Feng et al., 2024; Zeng et al., 2024). These models reframe the solution prediction as an iterative refinement process, decomposing the hard single-step task into a more manageable multi-step generation, which is better suited for navigating complex constrained spaces (Zeng et al., 2024). Despite these advances, prior approaches generally focus on partial distribution modeling on integer variables, leading to suboptimal guidance and inferior solutions. In contrast, our proposed Joint Continuous-Integer Flow for MILP is the first generative framework that explicitly models the joint distribution of both continuous and integer variables, directly addressing this critical gap in the literature.
@@ 3.1 MILP Definition and Graph Representation
Here, $\boldsymbol{w}$ is the objective function coefficient vector. $\boldsymbol{l}$ and $\boldsymbol{u}$ are the lower bounds and upper bounds for decision variables. $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ and $\boldsymbol{b}\in\mathbb{R}^{m}$ denote the coefficient matrix and the right-hand-side vector of linear constraints. Moreover, $K$ is a scalar integer defining the maximum feasible value of integer variables (i.e., $\boldsymbol{x}_{i}\in\{0,1,\dots,K\}$, $\forall i\in\mathcal{I}$). An MILP instance with $m$ linear constraints and $n$ variables can thus be represented by the tuple $\boldsymbol{M}=(\boldsymbol{A},\boldsymbol{b,l,u,w})$. A solution $\boldsymbol{x}$ is considered feasible if all constraints are satisfied. Note that, in this paper, we focus on MILP problems with bounded integer variables, as most real-world applications exhibit this characteristic. Specially, when $K=1$ such that $\boldsymbol{x}_{\mathcal{I}}\in\{0,1\}^{q}$, the problem reduces to Mixed Integer Binary Programming (MIBP), which is the most commonly studied subclass of MILP (Huang et al., 2024b; Liu et al., 2025).
Graph Representation for MILP. When applying deep learning to MILP, it is common to represent an MILP instance as a bipartite graph (Gasse et al., 2019a). The bipartite graph consists of two sets of nodes: one representing the $n$ decision variables and the other representing the $m$ constraints. Edges in this graph connect variables to the constraints in which they appear, effectively encoding the sparsity pattern of the constraint matrix. Building on this foundation, as shown in the top-left part of Figure 2, we employ a more granular tripartite graph structure to explicitly distinguish between variable types, which is crucial for our joint modeling approach. Specifically, an MILP instance is denoted as a graph $\boldsymbol{G}=(\mathcal{V}_{\mathcal{I}},\mathcal{V}_{\mathcal{C}},\mathcal{V}_{\text{con}},\mathcal{E})$, where the nodes are partitioned into three distinct sets:
@@
An edge is drawn between a variable node $j\in\mathcal{V}_{\mathcal{I}}\cup\mathcal{V}_{\mathcal{C}}$ and a constraint node $i\in\mathcal{V}_{con}$ if the variable $x_{j}$ has a non-zero coefficient in the $i$-th constraint. Based on the rich, relational MLIP graph, various graph neural networks (e.g., GCN (Kipf, 2016) or GAT (Veličković et al., 2017)) can be employed to extract the rich topological representations for later discriminative or generative learning processes. More details can be found in Appendix A and Appendix B.
@@ 4 FMIP: Joint Continuous-Integer Flow for MILP
In this section, we introduce FMIP, a generative framework that learns the joint distribution of integer and continuous variables to find high-quality solutions for MILP instances. As shown in Figure 2, FMIP is built on a time-dependent graph representation, a joint continuous-integer flow model, a holistic guidance mechanism for inference, and seamless integration with downstream solvers.
@@ 4.1 Time-Dependent Graph Representation
FMIP operates on a time-dependent solution graph, which captures both the static structure of an MILP instance and its dynamic state during the generative process. As defined in Section 3.1, the static structure is encoded in a tripartite graph $\boldsymbol{G}=(\mathcal{V}_{\mathcal{I}},\mathcal{V}_{\mathcal{C}},\mathcal{V}_{\text{con}},\mathcal{E})$.
@@ 4.3 Holistic Guidance Mechanism
Guidance Target Function. We define a target function $f(\boldsymbol{x})$ that captures the two primary goals of an MILP solution: minimizing the objective value and satisfying the constraints. The function is a weighted sum of the objective and a penalty for constraint violations:
@@
where $\boldsymbol{x}=(\boldsymbol{d},\boldsymbol{c})$ is the set of all variables, $\gamma$ is a balancing hyperparameter, and $\boldsymbol{A}_{i,*}$ is the $i$-th row of the constraint matrix. The model is guided to generate solutions that minimize this function. Following Lin et al. (2025), we implement guidance by modifying the update steps for the continuous and integer variables separately.
@@ 5.1 Experiment Settings
We evaluate FMIP on a comprehensive suite of eight MILP benchmarks. Five of them focus on classic combinatorial optimization problems: Combinatorial Auctions (CA) (Gasse et al., 2019a), Generalized Independent Set (GIS) (Colombi et al., 2017), Maximum Independent Set (MIS) (Gasse et al., 2019a), Fixed-Charge Multi-Commodity Network Flow (FCMNF) (Hewitt et al., 2010), and Set Covering (SC) (Gasse et al., 2019a). We also include two real-world MILP datasets with both binary and continuous variables from the NeurIPS ML4CO 2021 competition (Gasse et al., 2022): Load Balancing (LB) and Item Placement (IP). Finally, we adopt the standard MIPLIB2017 benchmark (MIPLIB) which is one of the most common standards for MILP solver evaluation (Gleixner et al., 2021a).
@@
FMIP and three baselines are all learning methods that are agnostic to backbone graph neural networks and downstream solvers. In later experiments, we employ four different backbone graph encoders, i.e., Tri-GCN (Gasse et al., 2019b), Bi-GCN (Gasse et al., 2019b), GAT (Brody et al., 2021), and ClusterGCN (Chiang et al., 2019), with Tri-GCN as the default choice. As for downstream solvers, we adopt Neural Diving (ND) (Nair et al., 2020), Predict-and-Search (PS) (Han et al., 2023), PMVB (Chen et al., 2025) and Apollo-MILP (Apollo) (Liu et al., 2025).
@@
Following previous works (Han et al., 2023; Liu et al., 2025), we report the objective value (OBJ) found by each method within a fixed time limit. To compare performance, we first determine the best-known solution (BKS), defined as the best objective value found across all methods, including a long run of Gurobi. We then calculate the absolute primal gap: $\text{GAP}=|\text{OBJ}-\text{BKS}|$. Since MIPLIB contains instances from diverse domains with very different optimal values, we report the relative primal gap for MIPLIB instead: $\text{Rel.GAP}=|\text{OBJ}-\text{BKS}|/(|\text{BKS}|+1)$. A lower absolute/relative primal gap indicates better performance.
@@
Due to the page limitation, we provide implementation details in Appendix C, including the backbone model architecture, training & inference hyperparameters of FMIP and baselines, as well as the downstream solver configuration.
@@ 5.2 Overall Performance
We choose Tri-GCN as the backbone graph encoder and evaluate FMIP against baselines based on four downstream solvers (ND, PS, PMVB, and Apollo), with respective time limits of 400, 600, 600, and 800 seconds. The time limit for Gurobi is set to 3600 seconds. The results are reported in Table 1. We observe that FMIP generally achieves the best performance among learning-based methods, substantially reducing the primal gap across all benchmarks and downstream solvers. On simpler benchmarks (i.e., CA, GIS) where multiple methods find the optimal solution, FMIP performs on par with the best baselines. However, on more challenging benchmarks (i.e., LB, IP), the benefits of our joint modeling and holistic guidance are clear, culminating in the significant performance gains in terms of both objective value and primal gap.
@@ 5.3 Compatibility Analysis
The downstream-solver compatibility of FMIP is already validated in Table 1 in Section 5.2, where FMIP consistently achieves the best performance across four different solvers. To further study the model compatibility of FMIP, we apply FMIP and baselines to four different backbone graph encoders on IP and LB benchmarks: Bi-GCN (Gasse et al., 2019b), Tri-GCN (Gasse et al., 2019b), GAT (Brody et al., 2021), and ClusterGCN (Chiang et al., 2019) We report the objective value (OBJ) metric in Table 2, where FMIP maintains its superior performance across all backbone graph encoders. This backbone independence demonstrates that FMIP acts as a powerful and general framework that can enhance a wide range of GNN-based models for various downstream solvers, highlighting the fundamental value of its joint distribution modeling on both continuous and integer variables.
Table 3: The inference time (s) per MILP instance of FMIP and baselines. We only compare the neural model inference here and exclude the time that downstream solvers take.
@@ 5.5 In-Depth Analysis
Ablation Study. We derive four variants of our proposed FMIP: (i) removing the objective guidance in Eq. 5 (w/o Feasibility Guidance), (ii) removing the objective guidance in Eq. 5 (w/o Objective Guidance), (iii) removing the entire holistic guidance mechanism (w/o Guidance), and (iv) reverting to an integer-only model by disabling continuous variable generation (w/o Continuous). As shown in Table 4, the full version of FMIP significantly outperforms all ablated variants, confirming that both the joint generation of all variables and the holistic guidance mechanism are critical to the performance gain. Notably, the “w/o Continuous” variant suffers from the worst performance, providing direct evidence that capturing the coupled relationship between integer and continuous variables is essential for high-quality solution prediction in MILP.
Impact of Sampling Steps. We analyze the effect of the number of sampling steps on the heuristic solution quality. Specifically, we feed the heuristic solutions from different sampling steps into the downstream solvers, and report the primal gap metric (i.e., GAP) in Figure 3. We can observe that, at the beginning, more sampling steps lead to better solutions as the model has more opportunities for refinement. However, the quality of heuristic solution starts to worsen when a plethora of steps are performed. We attribute this phenomenon to the trade-off between generative refinement and search diversity. An excessive number ofsampling steps might cause the guidance to make the predicted variable distribution overly sharp, which reduces the diversity of the solutions being explored and can prematurely trap the search in a local optimum. This insight reveals a nuanced interplay between generative refinement and search diversity, which is critical for optimizing performance.

### v3

# arXiv:2507.23390v3[math.OC] 28 Feb 2026

@@ 1 Introduction
Despite its versatility, solving a MILP instance remains a fundamental challenge due to its NP-hard property (Bertsimas and Tsitsiklis, 1997). Consequently, exact solvers often rely on heuristics to find high-quality solutions in a practical amount of time. This has motivated a surge of interest in using deep learning to predict powerful heuristics that can warm-start and guide the downstream solvers (Nair et al., 2020; Han et al., 2023; Liu et al., 2025; Chen et al., 2025). Among them, generative modeling has emerged as a particularly promising approach. Modern techniques like diffusion models and flow matching reframe the difficult one-step task of predicting a complete solution into an iterative refinement process (Lin et al., 2024b; Sun and Yang, 2023; Feng et al., 2024; Zeng et al., 2024). By learning to transform a simple noise distribution into a distribution of high-quality solutions, these models can effectively navigate the complex combinatorial search space, breaking down the challenge of satisfying intricate constraints into a series of more manageable steps.
@@
To this end, we propose Joint Continuous-Integer Flow for Mixed-Integer Linear Programming (FMIP), which is the first generative framework that models the joint distribution of both integer and continuous variables for MILP solutions. FMIP leverages a conditional flow matching process to progressively generate a complete solution, fully capturing the interdependence between all decision variables. This joint modeling approach unlocks a holistic guidance mechanism that can steer the generation process using complete, instance-wise feedback from both the objective function and constraint violations. We highlight that while the transition from ILP to MILP appears intuitive, joint modeling represents a critical but previously overlooked gap. Bridging this gap is crucial: 1) it enables us to shift from single-point prediction to generative distribution modeling, addressing the inherent “one-to-many” nature of MILP solutions (multiple optima), and 2) it enables the utilization of instance-wise optimization information, addressing a critical limitation of partial modeling schemes. FMIP is fully compatible with arbitrary backbone networks and downstream solvers, showing 41.34% relative improvement on average across eight standard MILP benchmarks over SOTA baselines. Our contributions are summarized as follows:
@@ 2 Related Works
Machine learning is widely used to accelerate Mixed-Integer Linear Programming (MILP) solvers. One major line of research focuses on enhancing the internal components of the Branch-and-Bound algorithm, such as learning policies for variable branching (Gasse et al., 2019; Khalil et al., 2016; Gupta et al., 2020; Zarpellon et al., 2021; Gupta et al., 2022; Scavuzzo et al., 2022; Lin et al., 2024a; Zhang et al., 2024) or cutting plane selection (Tang et al., 2020; Huang et al., 2022). Another closely related direction, which our work belongs to, aims to predict high-quality heuristic solutions to warm-start the optimization process, thereby reducing the search space and speeding up convergence (Nair et al., 2020; Han et al., 2023; Huang et al., 2024b; Liu et al., 2025; Chen et al., 2025; Zeng et al., 2024).
Most existing solution-prediction methods are discriminative in nature, employing models like Graph Neural Networks (GNNs) to predict the variable assignments in a single forward pass (Huang et al., 2024b; Liu et al., 2025; Hu et al., 2024; Huang et al., 2024a; Ye et al., 2024). However, predicting a complete, feasible, and high-quality solution in one step is exceptionally challenging due to the complex combinatorial structure of MILPs. This difficulty has motivated the use of generative models, such as diffusion models and flow matching, for MILP heuristics (Feng et al., 2024; Zeng et al., 2024). These models reframe the solution prediction as an iterative refinement process, decomposing the hard single-step task into a more manageable multi-step generation, which is better suited for navigating complex constrained spaces (Zeng et al., 2024). However, we identify a fundamental limitation in these state-of-the-art approaches: prior approaches generally focus on partial distribution modeling on integer variables, which inherently precludes capturing the critical interdependence between all decision variables, leading to suboptimal guidance and inferior solutions. In contrast, our proposed Joint Continuous-Integer Flow for MILP is the first generative framework that explicitly models the joint distribution of both continuous and integer variables, directly addressing this critical gap in the literature.
@@ 3.1 MILP Definition and Graph Representation
Here, $\boldsymbol{w}$ is the objective function coefficient vector. $\boldsymbol{l}$ and $\boldsymbol{u}$ are the lower bounds and upper bounds for decision variables. $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ and $\boldsymbol{b}\in\mathbb{R}^{m}$ denote the coefficient matrix and the right-hand-side vector of linear constraints. Moreover, $K$ is a scalar integer defining the maximum feasible value of integer variables (i.e., $\boldsymbol{x}_{i}\in\{0,1,\dots,K\}$, $\forall i\in\mathcal{I}$). An MILP instance with $m$ linear constraints and $n$ variables can thus be represented by the tuple $\boldsymbol{M}=(\boldsymbol{A},\boldsymbol{b,l,u,w})$. A solution $\boldsymbol{x}$ is considered feasible if all constraints are satisfied. Note that, in this paper, we focus on MILP problems with bounded integer variables, as most real-world applications exhibit this characteristic. Specially, when $K=1$ such that $\boldsymbol{x}_{\mathcal{I}}\in\{0,1\}^{q}$, the problem reduces to Mixed Integer Binary Programming (MIBP), which is the most commonly studied subclass of MILP (Huang et al., 2024b; Liu et al., 2025). Our FMIP framework can naturally extend to unbounded integer variables by incorporating the bit-wise prediction strategy employed in prior studies (Nair et al., 2020; Han et al., 2023). Alternatively, we can continuously relax a subset of these variables to predict their continuous embeddings, which are subsequently discretized via rounding or projection to obtain valid integer assignments.
Graph Representation for MILP. When applying deep learning to MILP, it is common to represent an MILP instance as a bipartite graph (Gasse et al., 2019). The bipartite graph consists of two sets of nodes: one representing the $n$ decision variables and the other representing the $m$ constraints. Edges in this graph connect variables to the constraints in which they appear, effectively encoding the sparsity pattern of the constraint matrix. Building on this foundation, as shown in the top-left part of Figure 2, we maintain the fundamental bipartite topology but explicitly distinguish between integer and continuous variables as distinct node types. Specifically, an MILP instance is denoted as a graph $\boldsymbol{G}=(\mathcal{V}_{\mathcal{I}},\mathcal{V}_{\mathcal{C}},\mathcal{V}_{\text{con}},\mathcal{E})$, where the nodes are partitioned into three distinct sets:
@@
An edge is drawn between a variable node $j\in\mathcal{V}_{\mathcal{I}}\cup\mathcal{V}_{\mathcal{C}}$ and a constraint node $i\in\mathcal{V}_{con}$ if the variable $x_{j}$ has a non-zero coefficient in the $i$-th constraint. Based on the rich, relational MLIP graph, various graph neural networks (e.g., GCN (Kipf, 2016) or GAT (Veličković et al., 2017)) can be employed to extract the rich topological representations for later discriminative or generative learning processes. More details can be found in Appendix B and Appendix C.
@@ 4 FMIP: Joint Continuous-Integer Flow for MILP
In this section, we introduce FMIP, a generative framework that learns the joint distribution of integer and continuous variables to find high-quality solutions for MILP instances. This unified modeling is the cornerstone of our approach, explicitly designed to capture variable coupling and enable the utilization of instance-wise optimization information, which is inaccessible to partial modeling. As shown in Figure 2, FMIP is built on a time-dependent graph representation, a joint continuous-integer flow model, a holistic guidance mechanism for inference, and seamless integration with downstream solvers. Notably, our FMIP framework is independent of the choice of generative model. We employ flow matching in this work due to its conceptual simplicity and training stability.
@@ 4.1 Time-Dependent Graph Representation
FMIP operates on a time-dependent solution graph, which captures both the static structure of an MILP instance and its dynamic state during the generative process. As defined in Section 3.1, the static structure is encoded in a graph $\boldsymbol{G}=(\mathcal{V}_{\mathcal{I}},\mathcal{V}_{\mathcal{C}},\mathcal{V}_{\text{con}},\mathcal{E})$.
@@ 4.3 Holistic Guidance Mechanism
Guidance Target Function. We define a target function $f(\boldsymbol{x})$ that captures the two primary goals of an MILP solution: minimizing the objective value and the constraint violation. The function is a weighted sum of the objective and a penalty for constraint violations:
@@
where $\boldsymbol{x}=(\boldsymbol{d},\boldsymbol{c})$ is the set of all variables, $\gamma$ is the weight of constraint-violation term to balance two guidance parts, and $\boldsymbol{A}_{i,*}$ is the $i$-th row of the constraint matrix. The model is guided to generate solutions that minimize this function. Following Lin et al. (2025), we implement guidance by modifying the update steps for the continuous and integer variables separately.
@@ 5.1 Experiment Settings
We evaluate FMIP on a comprehensive suite of eight MILP benchmarks. Five of them focus on classic combinatorial optimization problems: Combinatorial Auctions (CA) (Gasse et al., 2019), Generalized Independent Set (GIS) (Colombi et al., 2017), Maximum Independent Set (MIS) (Gasse et al., 2019), Fixed-Charge Multi-Commodity Network Flow (FCMNF) (Hewitt et al., 2010), and Set Covering (SC) (Gasse et al., 2019). We also include two real-world MILP datasets with both binary and continuous variables from the NeurIPS ML4CO 2021 competition (Gasse et al., 2022): Load Balancing (LB) and Item Placement (IP). Finally, we adopt the standard MIPLIB2017 benchmark (MIPLIB) which is one of the most common standards for MILP solver evaluation (Gleixner et al., 2021a).
@@
FMIP and three baselines are all learning methods that are agnostic to backbone graph neural networks and downstream solvers. In later experiments, we employ four different backbone graph encoders, i.e., Tri-GCN (Gasse et al., 2019), Bi-GCN (Gasse et al., 2019), GAT (Brody et al., 2021), and ClusterGCN (Chiang et al., 2019), with Tri-GCN as the default choice. Note that Tri-GCN and Bi-GCN both operate on a bipartite topology; the former is distinguished by handling integer and continuous variables as heterogeneous node types with distinct parameters. As for downstream solvers, we adopt Neural Diving (ND) (Nair et al., 2020), Predict-and-Search (PS) (Han et al., 2023), PMVB (Chen et al., 2025) and Apollo-MILP (Apollo) (Liu et al., 2025).
@@
Following previous works (Han et al., 2023; Liu et al., 2025), we report the objective value (OBJ) found by each method within a fixed time limit. To compare performance, we first determine the best-known solution (BKS), defined as the best objective value found across all methods, including a long run of Gurobi. We then calculate the absolute primal gap: $\text{GAP}=|\text{OBJ}-\text{BKS}|$. Since MIPLIB contains instances from diverse domains with very different optimal values, we report the relative primal gap for MIPLIB instead: $\text{Rel.GAP}=|\text{OBJ}-\text{BKS}|/(|\text{BKS}|+1)$. A lower absolute/relative primal gap indicates better performance. Notably, we only report the Primal Gap instead of the Optimal Gap because FMIP acts as a primal heuristic. Its objective is to quickly identify high-quality feasible solutions (improving the upper bound), rather than tightening the dual bound/optimality gap. This is the standard metric for evaluating heuristic improvements.
@@
Due to the page limitation, we provide implementation details in Appendix D, including the backbone model architecture, training & inference hyperparameters of FMIP and baselines, as well as the downstream solver configuration.
@@ 5.2 Overall Performance
We choose Tri-GCN as the backbone graph encoder and evaluate FMIP against baselines based on four downstream solvers (ND, PS, PMVB, and Apollo), with respective time limits of 400, 600, 600, and 800 seconds. The time limit for Gurobi is set to 3600 seconds. The results are reported in Table 1. We observe that FMIP generally achieves the best performance among learning-based methods, substantially reducing the primal gap across all benchmarks and downstream solvers. On simpler benchmarks (i.e., CA, GIS) where multiple methods find the optimal solution, FMIP performs on par with the best baselines. However, on more challenging benchmarks (i.e., LB, IP), the benefits of our joint modeling and holistic guidance are clear, culminating in the significant performance gains in terms of both objective value and primal gap. Overall, our method brings a 41.34% improvement compared with the best machine learning baseline, calculated as the mean increase over all datasets and downstream solvers.
@@ 5.3 Compatibility Analysis
The downstream-solver compatibility of FMIP is already validated in Table 1 in Section 5.2, where FMIP consistently achieves the best performance across four different solvers. To further study the model compatibility of FMIP, we apply FMIP and baselines to four different backbone graph encoders on IP and LB benchmarks: Bi-GCN (Gasse et al., 2019), Tri-GCN (Gasse et al., 2019), GAT (Brody et al., 2021), and ClusterGCN (Chiang et al., 2019) We report the objective value (OBJ) metric in Table 2, where FMIP maintains its superior performance across all backbone graph encoders. This backbone independence demonstrates that FMIP acts as a powerful and general framework that can enhance a wide range of GNN-based models for various downstream solvers, highlighting the fundamental value of its joint distribution modeling on both continuous and integer variables.
Table 3: The inference time (s) per MILP instance of FMIP and baselines. We only compare the neural model inference here, including the computation of guidance, and exclude the time that downstream solvers take.
@@ 5.5 In-Depth Analysis
Ablation Study. We derive four variants of our proposed FMIP: (i) removing the objective guidance in Eq. 5 (w/o Feasibility Guidance), (ii) removing the objective guidance in Eq. 5 (w/o Objective Guidance), (iii) removing the entire holistic guidance mechanism(w/o Guidance), and (iv) reverting to an integer-only model by disabling continuous variable generation (w/o Continuous), which inherently prevents the application of holistic guidance. As shown in Table 4, the full version of FMIP significantly outperforms all ablated variants, confirming that both the joint generation of all variables and the holistic guidance mechanism are critical to the performance gain. Notably, the “w/o Continuous” variant suffers from the worst performance, providing direct evidence that capturing the coupled relationship between integer and continuous variables is essential for high-quality solution prediction in MILP.
Impact of Sampling Steps. We analyze the effect of the number of sampling steps on the heuristic solution quality. Specifically, we feed the heuristic solutions from different sampling steps into the downstream solvers, and report the primal gap metric (i.e., GAP) in Figure 3. We can observe that, at the beginning, more sampling steps lead to better solutions as the model has more opportunities for refinement. However, the quality of heuristic solution starts to worsen when a plethora of steps are performed. We attribute this phenomenon to the trade-off between generative refinement and search diversity. An excessive number of sampling steps might cause the guidance to make the predicted variable distribution overly sharp, which reduces the diversity of the solutions being explored and can prematurely trap the search in a local optimum. Alternatively, this degradation may stem from the balance between objective value and feasibility in the target function 5, which is discussed in the subsequent part.
@@
Adaptive Guidance Balancing. We investigate the potential imbalance between objective and feasibility guidance terms during iterative generation. We hypothesize that a fixed coefficient $\gamma$ (defaulting to 50) leads to performance degradation over longer sampling trajectories. To address this, we introduce an adaptive balancing strategy that dynamically adjusts $\gamma$ at each generation step. Specifically, we compare the magnitude of the objective guidance term with that of the constraint-violation term: if the objective term dominates, we scale up $\gamma$ by a factor $\alpha$ (i.e., $\gamma\leftarrow\gamma\times\alpha$) to enforce stronger constraint satisfaction; conversely, if the constraint term prevails, we scale down $\gamma$ (i.e., $\gamma\leftarrow\gamma/\alpha$). This adjustment is cumulative across sampling steps, allowing the model to dynamically recalibrate the guidance focus between optimality and feasibility. We evaluate this strategy on the Item Placement dataset using the PS solver, varying $\alpha$ and the number of steps. As shown in Table 5, with a fixed $\gamma$, the objective value worsens significantly as steps increase (e.g., from 11.79 at Step-12 to 13.25 at Step-32). However, the adaptive strategy ($\alpha>1$) effectively mitigates this degradation, maintaining competitive performance even at deeper generation steps (e.g., 11.92 at Step-32 with $\alpha=2.0$). These results validate that dynamically balancing guidance is crucial for stabilizing long-trajectory generation.
Table 5: Impact of Adaptive Balancing Strategy on Objective Value (Lower is Better) across different sampling steps on the IP dataset.
Strategy Step-12 Step-24 Step-32
Fixed $\gamma$ 11.79 12.39 13.25
Adaptive ($\alpha=1.1$) 11.93 12.35 12.05
Adaptive ($\alpha=1.5$) 11.99 11.85 12.68
Adaptive ($\alpha=2.0$) 12.19 11.95 11.92